Mr Sprayberry is running late to school. The speed limit is 40 miles per hour, and he is stopped at a stop light. After the light turns green, he accelarates at 5.2 m/s for 19 seconds. Is he speeding a the end of his accelaration? What would his acceleration be for those 90 seconds in order to reach exactly 40 miles per hour?
Answer #1

Acceleration= 5.2 m/s​​​​​​2

time taken = 19 seconds

Using V = U + at

V = 0 + (5.2)(19)

V= 98.8 m/s

V = 98.8 × 2.237 mph

V = 221 mph

Therefore, yes he is over speeding as the limit is only 40 mph but he is traveling at 221 mph .

Now , t = 90 second

V = 40 mph

V = 40/2.237 m/s

V = 17.88 m/s

Using, V = U +at

17.88 = 0 + (a)(90)

a = 0.1986 m/s 2

Therefore he should accelerate at the rate of 0.1986 m/s 2 .

If time was 19 seconds, instead of 90 seconds then acceleration would have been = 0.94 m/s 2